OpenStudy (anonymous):

The formula below gives V, the volume of a cylinder with radius r and height 14. V(r) = 14(pie)r2 What is the value of V(3)? A. 84 B. 23 C. 112 D. 126 There's a pie symbol between the 14 and r. its not in parentheses.

OpenStudy (phi):

btw, you mean pi (greek P). pie is something you eat. v(3) is short-hand for: find the definition of v( r ) and everywhere you see r in the formula, replace the r with 3 then simplify can you do that ?

OpenStudy (anonymous):

would it be 126(pi) ?

OpenStudy (phi):

it would be \[ V(3) = 14\pi 3^2 \\ = 14\cdot 9 \pi \\ = 126 \pi\]

OpenStudy (phi):

so yes

OpenStudy (anonymous):

Okay thanks! that's how I did it.